# A proposal for removing $\pi N$-state contamination from the nucleon induced pseudoscalar form factor in lattice QCD

### v1

@@ 1 Introduction
with $q=p^{\prime}-p$, a specific relation, known as the Goldberger-Treiman (GT) relation [1], is derived between the axial-vector coupling defined by the axial-vector ($F_{A}$) form factor at $q^{2}=0$ and the residue of the pion-pole structure in the induced pseudo-scalar ($F_{P}$) form factor. Instead of PCAC, the axial Ward-Takahashi identity, $\partial_{\alpha}A_{\alpha}(x)=2mP(x)$, leads to the generalized GT relation [2, 3]:
@@
Although the axial-structure of the nucleon has been studied extensively in lattice QCD at the physical point, no results have been obtained that satisfy the generalized GT relation well or give better accuracy than the PPD model. Indeed, in our previous works [5, 6], both the $F_{P}$ and $G_{P}$ form factors are significantly underestimated in the low-$q^{2}$ region compared to the PPD model. This is simply due to strong $\pi N$ excited-state contamination. In this work, we propose a simple method to remove the $\pi N$-state contamination from the $F_{P}$ and $G_{P}$ form factors, and then evaluate the induced pseudoscalar charge $g_{P}^{\ast}=m_{\mu}F_{A}(0.88m_{\mu}^{2})$ and the pion-nucleon coupling $g_{\pi NN}=\lim_{q^{2}\rightarrow\infty}(q^{2}+m_{\pi}^{2})\frac{F_{P}(q^{2})}{2F_{\pi}}$ from existing data in a new analysis.
@@ Acknowledgments
We would like to thank members of the PACS collaboration for useful discussions. K.-I. I. is supported in part by MEXT as “Feasibility studies for the next-generation computing infrastructure". Numerical calculations in this work were performed on Oakforest-PACS in Joint Center for Advanced High Performance Computing (JCAHPC) and Cygnus and Pegasus in Center for Computational Sciences at University of Tsukuba under Multidisciplinary Cooperative Research Program of Center for Computational Sciences, University of Tsukuba, and Wisteria/BDEC-01 in the Information Technology Center, the University of Tokyo. This research also used computational resources of the K computer (Project ID: hp1810126) and the Supercomputer Fugaku (Project ID: hp20018, hp210088, hp230007, hp2320199, hp240207) provided by RIKEN Center for Computational Science (R-CCS), as well as Oakforest-PACS (Project ID: hp170022, hp180051, hp180072, hp190025, hp190081, hp200062), Wisteria/BDEC-01 Odyssey (Project ID: hp220050) provided by the Information Technology Center of the University of Tokyo / JCAHPC. The calculation employed OpenQCD system(http://luscher.web.cern.ch/luscher/openQCD/). This work is supported by the JLDG constructed over the SINET5 of NII, This work was also supported in part by Grants-in-Aid for Scientific Research from the Ministry of Education, Culture, Sports, Science and Technology (Nos. 18K03605, 19H01892, 22K03612, 23H01195, 23K03428, 23K25891) and MEXT as “Program for Promoting Researches on the Supercomputer Fugaku” (Search for physics beyond the standard model using large-scale lattice QCD simulation and development of AI technology toward next-generation lattice QCD; Grant Number JPMXP1020230409).

### v2

@@ 1 Introduction
with $q=p-p^{\prime}$, a specific relation, known as the Goldberger-Treiman (GT) relation [1], is derived between the axial-vector coupling defined by the axial-vector ($F_{A}$) form factor at $q^{2}=0$ and the residue of the pion-pole structure in the induced pseudo-scalar ($F_{P}$) form factor. Instead of PCAC, the axial Ward-Takahashi identity, $\partial_{\alpha}A_{\alpha}(x)=2mP(x)$, leads to the generalized GT relation [2, 3]:
@@
Although the axial-structure of the nucleon has been studied extensively in lattice QCD at the physical point, no results have been obtained that satisfy the generalized GT relation well or give better accuracy than the PPD model. Indeed, in our previous works [5, 6], both the $F_{P}$ and $G_{P}$ form factors are significantly underestimated in the low-$q^{2}$ region compared to the PPD model. This is simply due to strong $\pi N$ excited-state contamination. In this work, we propose a simple method to remove the $\pi N$-state contamination from the $F_{P}$ and $G_{P}$ form factors, and then evaluate the induced pseudoscalar charge $g_{P}^{\ast}=m_{\mu}F_{P}(0.88m_{\mu}^{2})$ and the pion-nucleon coupling $g_{\pi NN}=\lim_{q^{2}\rightarrow\infty}(q^{2}+m_{\pi}^{2})\frac{F_{P}(q^{2})}{2F_{\pi}}$ from existing data in a new analysis.
@@ Acknowledgments
We would like to thank members of the PACS collaboration for useful discussions. K.-I. I. is supported in part by MEXT as “Feasibility studies for the next-generation computing infrastructure". Numerical calculations in this work were performed on Oakforest-PACS in Joint Center for Advanced High Performance Computing (JCAHPC) and Cygnus and Pegasus in Center for Computational Sciences at University of Tsukuba under Multidisciplinary Cooperative Research Program of Center for Computational Sciences, University of Tsukuba, and Wisteria/BDEC-01 in the Information Technology Center, the University of Tokyo. This research also used computational resources of the K computer (Project ID: hp180126) and the Supercomputer Fugaku (Project ID: hp20018, hp210088, hp230007, hp230199, hp240028, hp240207) provided by RIKEN Center for Computational Science (R-CCS), as well as Oakforest-PACS (Project ID: hp170022, hp180051, hp180072, hp190025, hp190081, hp200062), Wisteria/BDEC-01 Odyssey (Project ID: hp220050) provided by the Information Technology Center of the University of Tokyo / JCAHPC. The calculation employed OpenQCD system(http://luscher.web.cern.ch/luscher/openQCD/). This work is supported by the JLDG constructed over the SINET5 of NII, This work was also supported in part by Grants-in-Aid for Scientific Research from the Ministry of Education, Culture, Sports, Science and Technology (Nos. 18K03605, 19H01892, 22K03612, 23H01195, 23K03428, 23K25891) and MEXT as “Program for Promoting Researches on the Supercomputer Fugaku” (Search for physics beyond the standard model using large-scale lattice QCD simulation and development of AI technology toward next-generation lattice QCD; Grant Number JPMXP1020230409).